A is a constant such that the graph of the equation Ax-3y=6 passes through the point (1,-3). Find A
Found 2 solutions by Theo, josgarithmetic:
Answer by Theo(13342)

the equation is Ax - 3y = 6
it passed through the point (1,-3).
replace x with 1 and y with -3 to get:
A * 1 - 3 * (-3) = 6
simplify to get:
A + 9 = 6
subtract 9 from both sides of the equation to get:
A = 6 - 9
simplify to get:
A = -3.

when A = -3, the equation of Ax - 3y = 6 becomes -3x - 3y = 6
when x = 1 and y = -3, the equation becomes:
-3 * 1 - 3 * (-3) = 6
simplify to get:
-3 + 9 = 6
simplify to get:
6 = 6
this confirms the value of A is good.

using the desmos.com calculator, this equation can be graphed as is:
here's what the graph looks like.



as you can see, it's a straight line that passes through the point (1,-3).

your solution is that A = -3.

Answer by josgarithmetic(39617)